STATEMENT OF CHANGES IN EQUITY  

PRACTICE SET 29 

  

  1. The Retrieval Company has the following changes in equity items in year-end 2020.  
  2. Kindly prepare the SCE of the company which is owned by Ms. Charlott 

  

Balance on January 1, 2020= ₱ 550,000 

  

March 31: The entity earned a total of ₱ 250,000 for its first quarter of operations 

June 30: The owner withdraws money worth ₱ 300,000 

July 28: Ms. Charlott, the owner, invests a total of ₱ 920,000 worth of equipment in the entity. 

September 22: The entity earned a total of ₱ 450,000  

December 31: Total expenses paid at year-end amounts to ₱ 380,000 

The owner withdraws 500,000 from the business 

  

Balance on December 31, 2020=? 

  

  


1
Expert's answer
2021-10-13T10:40:03-0400

Owner's withdrawal = ₱ 300,000 + 500,000 = ₱ 800,000

Total earnings = ₱ 250,000 + ₱ 450,000 = ₱ 700,000

Total Expenses = 380,000

SCE= 550,000 + 920,000 +(700,000 - 380,000) - 800,000 = 990,000

The Retrieval Company

Statement of Changes in Equity

For the year ended 31st December 2020

Capital opening balance Jan1,2020. ₱ 550,000

Add: Investment (Equipment) ₱ 920,000

Add: Net Income (Total earnings - Total Expenses). ₱ 320,000

Less: Total Withdrawal. ₱ 800,000

Capital Closing Balance Dec 31,2020 = ₱ 990,000